# Magnetosphere Evolution and Precursor-Driven Electromagnetic Signals in Merging Binary Neutron Stars

### v1

@@ 2.3 Specification of Orbital Inspiral
In this work, we consider the gravitational radiation-driven inspiral dynamics of the BNS system. With regard to these dynamics, we neglect any backreaction from the evolving magnetosphere, and we treat the NSs as point masses, thereby ignoring tidal distortions, dynamical tides (T. Hinderer et al., 2016) or energy transfers related to NS oscillation modes (D. Lai, 1994; A. Reisenegger & P. Goldreich, 1994; K. D. Kokkotas & G. Schafer, 1995), which may excited during inspirals.
In particular, for the low-eccentricity inspiral orbits of the two NSs, with masses $M_{1}$ and $M_{2}$, we follow the description in P. C. Peters & J. Mathews (1963); P. C. Peters (1964); L. D. Landau & E. M. Lifschits (1975) for the simplest case of point masses and the GW mass quadrupole formula. The length and frequency scales are normalized using a characteristic gravitational radius [^4] This corresponds to the horizon of the maximally spinning Kerr black hole with the total mass of the system., $R_{g}=G(M_{1}+M_{2})/c^{2}$ where $G$ is Newton’s gravitational constant and $c$ the speed of light, and the Keplerian frequency at $R_{g}$, $\Omega_{g}=[G(M_{1}+M_{2})/R_{g}^{3}]^{1/2}$, respectively. The separation $d$ between the two NSs is expressed in terms of the orbital angular frequency $\Omega$ as
@@
where $\Omega_{i}$ is the initial orbital frequency and
@@
In our simulations, the two NSs are treated as solid conducting bodies occupying finite and fixed volumes throughout the entire evolution. Consequently, the inspiral is considered to end when the surfaces of the two stars touch, which occurs at an earlier time, $t_{\rm Final}$, than the formal merger time predicted for point masses from the GW mass quadrupole formula. This corresponds to a separation of $d_{\rm Final}=2r_{\star}$, where $r_{\star}$ is the radius of each NS. The time $t_{\rm Final}$ reads (M. V. Medvedev & A. Loeb, 2013):
@@
where for the adopted values $r_{\star}=12\rm\,km$ and $M_{1}=M_{2}=1.4M_{\odot}$, this corresponds to a duration of $t_{0}~{}-~{}t_{\rm Final}~{}\sim~{}1.2\rm\,ms$. Hereafter, we define $\tau=t-t_{0}$ as the simulation time indicator, which is always negative, representing the time remaining until the merger. Although $\tau$ increases algebraically as the system evolves toward merger, its absolute value decreases over time.
@@ 2.4 Magnetospheric Evolution and Numerical Implementation
As the conductivity $\sigma$ varies from 00 to $\infty$, the magnetosphere structure spans the entire spectrum of solutions, from vacuum to FF. For practical purposes, we adopt a relatively high value of conductivity, $\sigma=6\Omega_{\rm i}$, which brings the system close to the FF regime. At the end of each time step, any remaining parallel electric field component is subtracted to further enforce the ideal FF condition.
During the evolution, however, there is a possibility for regions to develop where $E>B$ (A. Spitkovsky, 2006; C. Kalapotharakos & I. Contopoulos, 2009). Such regions are inconsistent with the FF regime due to the presence of accelerating electric field components. To prevent this, we enforce the condition $E\leq B$ throughout the simulation domain at the end of each time step. In realistic FF conditions, the plasma would be redistributed in regions where $E$ exceeds $B$ to neutralize the corresponding accelerating electric field component, restoring consistency with the FF approximation.
Our FDTD code integrates the Maxwell equations 89, tracking the inspiral of the two NSs by enforcing field values inside them that correspond to their respective magnetic moments, with the electric fields determined by Eq. (3). Compared to simulations of an isolated spinning NS, the merging BNS system proved more susceptible to the growth of oscillations (spurious ripples) near regions with steep gradients or discontinuities, such as current sheets. To address this, we implemented a 3rd-order Total Variation Diminishing (TVD) Runge-Kutta scheme (C.-W. Shu & S. Osher, 1988), replacing the standard 3rd-order Runge-Kutta method used previously. This TVD scheme is well-known for its enhanced numerical stability, particularly in handling sharp transitions and avoiding oscillatory artifacts.
In the case of perfect conductors, as the stars are modeled here, the boundary conditions require the continuity of the tangential components of the electric field and the normal components of the magnetic field. Handling spherical boundaries within a Cartesian grid poses significant computational challenges. This issue arises even in the simpler case of an isolated spinning NS. However, in the BNS system, the motion of the stellar boundaries aggravates the problem, often spouting unphysical non-zero values of $\nabla\cdot\mathbf{B}$ near the stellar boundary as the system evolves.
@@
Figure 3: The total Poynting flux that crosses a surface of radius $13\,r_{\star}$ during systems evolution starting from the initial separation of $3.3\,r_{\star}$ (red line) and $8\,r_{\star}$ (black line). Both simulations show similar flux evolution in the case of $\alpha_{1}=0^{\circ},\,\alpha_{2}=45^{\circ}$ and $\lambda=-1$ (
@@
To address this, we implement the diffusive method of M. Zhang & X. Feng (2016), which adds an additional term to Eq. 9 to eliminate any residual divergence of the magnetic field. In practice, at the end of each time step, we apply the following scheme
@@
The parameter $\eta$ has to satisfy (see M. Zhang & X. Feng, 2016, and references therein),
@@
The evolution of the binary system is tracked from an initial separation, $d=3.3r_{\star}$, until the stars contact, which from Equation 7 corresponds to $t_{\rm Final}=7.7\rm\,ms$ with a timestep of $dt=0.04\,r_{\star}/c$ and an initial orbital frequency ranging from $\Omega_{\rm i}\sim 2451\rm\,\,rad/s\,\,(390\,\rm Hz)$ to $\Omega_{\rm Final}\sim 5195\rm\,\,rad/s\,\,(827\,\rm Hz)$. The grid consists of $512$ grid points in each direction, covering the entire computational domain with an edge length of $25.6r_{\star}$ centered on the binary system. However, the active computational domain, excluding the PML region, has an edge length of $22.3r_{\star}$. This region, in practice, is where the simulation’s magnetospheric evolution is calculated.
Resolution studies were also conducted. For a selected subset of configurations, we performed simulations at higher spatial resolution and smaller time steps, utilizing 1024 grid points per dimension, effectively doubling the resolution, or with computational domains twice the size of the standard setup. The results presented in §3 were found to be robust, showing no significant variation with enhanced resolution and box size.
@@ 3 Results
A series of simulations were run to survey different unique initial configurations for the magnetic fields of the two stars. We explored $\lambda=\{\pm 1,\pm 10\}$, $\alpha_{1},\alpha_{2}=\{0^{\circ},45^{\circ},90^{\circ}\}$ and $\Delta\Phi=\{0^{\circ},45^{\circ},90^{\circ}\}$. Additionally, as changes in the $\Delta\Phi$ parameter only slightly affect the outflow patterns, for $\Delta\Phi=0$ further cases were explored with $\lambda=\{\pm\sqrt{10},\pm 100\}$ and the same range for the inclination angles. Lastly, to validate the specific results presented below (see section  3.1.5), we also performed additional simulations for cases with $\lambda=\{\pm\sqrt{10^{3}},\pm\sqrt{10^{5}},\pm 10^{3}\}$, $\alpha_{1}=\{0^{\circ},45^{\circ}\}$, $\alpha_{2}=0$ and $\Delta\Phi=0$. This exploration resulted in $24132=(3^{3}+3^{2}+3)2^{2}-$ 24 unique [^5] When $\alpha_{2}=0$, the configuration is the same for every $\Delta\Phi$, thus those cases have been excluded simulations.
@@
, left two columns) and anti-aligned (
@@
, right two columns) magnetic moments on the $z-x$ plane in the co-rotating frame. Black streamlines indicate the magnetic field direction on the poloidal plane, while the color denotes the normalized magnitude of the magnetic field component perpendicular to it (first and third column) and the magnitude of the magnetic field (second and fourth column). Dashed vertical lines indicate the characteristic distance $c/\Omega$.
@@ 3.1 Electromagnetic Fluxes, Field Structures, and Emission Anisotropy
For a single pulsar with a plasma-filled magnetosphere, as it rotates around its spin axis with an angular frequency $\omega$, energy is lost to low-frequency EM radiation that is mostly sourced from the open magnetic field lines (I. Contopoulos et al., 1999; A. Spitkovsky, 2006; N. Bucciantini et al., 2006; C. Kalapotharakos & I. Contopoulos, 2009; A. Tchekhovskoy et al., 2016). Calculating the magnetic flux along the open magnetic field lines $\Phi_{\rm open}$, one may compute an approximate luminosity as $L\approx\omega^{2}\Phi_{\rm open}^{2}/(6\pi^{2}c)$ for direct comparison with the integrated Poynting flux crossing a surface at a large distance.
[…]
@@
), presented at two snapshots separated by half orbital rotation. Bottom row: A configuration $\alpha_{1}=90^{\circ},\alpha_{2}=90^{\circ},\Delta\Phi=0^{\circ}$ and $\lambda=-1$ (
@@
In particular, the outgoing EM power can be calculated by integrating the radial component of the Poynting vector, $\mathbf{S}=c\mathbf{E}\times\mathbf{B}/(4\pi)$, over a spherical surface of radius $r_{F}$. Numerical FF simulations (A. Spitkovsky, 2006), later validated by magnetohydrodynamics and particle-in-cell (PIC) simulations within the FF regime, have shown that the luminosity of a single isolated pulsar with a magnetic inclination angle $\alpha$ follows an approximate relation
@@
The Poynting flux distribution crossing a spherical surface at a given time (Fig. 2, left panel) rotates in sync with the pulsar’s rotation. By integrating this luminosity over a full rotation period, the average luminosity distribution can be determined (Fig. 2, right). An observer at infinity may be pointed at a specific set of azimuthal angle $\phi$ and polar angle $\zeta$, i.e., the so-called observer angle relative to the spin axis. Notably, the average luminosity observed over a complete rotation depends exclusively on $\zeta$.
@@
[…]
The binary system, starting from a large initial separation, evolves in a low orbital frequency regime as dictated by Eq. (4), resulting in a gradual evolution toward merger. For a synchronized system (where the stars have the same spin and orbital frequencies) in this regime, the light cylinder is mathematically well-defined at a distance $R_{\rm LC}=c/\Omega$. In this configuration, the magnetosphere structure corotates with the system (neglecting the radial inspiral motion), and the Poynting flux distribution forms a fixed, rotating pattern. However, in this study, the stars are effectively irrotational, and no strictly defined single light cylinder exists. Instead, the length scale $c/\Omega$ can be interpreted as an instantaneous light cylinder associated with the evolving field configuration and orbital frequency $\Omega$. Yet, $c/\Omega$ is a crucial length scale for the magnetosphere’s evolution, and the active computational domain size is chosen so that this length scale is always well within it.
For the initial orbital frequency $\Omega_{i}$, we have $c/\Omega_{i}=10.2r_{\star}$, while at the final time $t=t_{\text{final}}$, it decreases to $c/\Omega_{f}=4.8r_{\star}$ owing to the inspiral.
@@
[…]
3.1.1 Aligned and Anti-aligned Magnetic Moments
[…]
The primary distinction between these two cases lies in the behavior of the magnetic field lines in the region around the two stars. If each star were isolated, these magnetic field lines would belong to either the closed or open magnetic field regions of their respective magnetospheres. In the presence of both stars, these lines do not maintain the same behavior. In the anti-aligned case, some magnetic field lines connect the two stars directly. In contrast, in the aligned case, these field lines are bent out of the poloidal plane in the azimuthal direction, closing within the opposite hemisphere of the star they originate from. This behavior is more evident in a 3D representation of the magnetic field lines, as shown in the top row of Figure 5. The left column shows the magnetic field lines from the equatorial plane, while the right column provides a top-down view. Due to the symmetry, the field lines of each star would have the same directions, and thus, no line is allowed to connect the two stars. Furthermore, as each star’s magnetic field lines are repelled from the region dominated by the other star’s magnetosphere, they become compressed in the leading direction of motion. In contrast, comparatively fewer field lines populate the trailing side toward the companion. Due to retardation effects, the field lines are consistently swept back, resulting in an “S”-shaped morphology when viewed from above, i.e., along the $z$-axis (see left panel of the top row in Fig. 5). The motion of the stars also results in a constant opening and closing of field lines, producing a stable flow of EM flux along the open field lines that increases following the increase of the fields as the separation decreases.

### v2

@@ 2.3 Specification of Orbital Inspiral
In this work, we consider the gravitational radiation-driven inspiral dynamics of the BNS system. With regard to these dynamics, we neglect any backreaction from the evolving magnetosphere, and we treat the NSs as point masses, thereby ignoring tidal distortions, dynamical tides (T. Hinderer et al., 2016), or energy transfers related to NS oscillation modes (D. Lai, 1994; A. Reisenegger & P. Goldreich, 1994; K. D. Kokkotas & G. Schafer, 1995), which may be excited during inspirals.
In particular, for the low-eccentricity inspiral orbits of the two NSs, with masses $M_{1}$ and $M_{2}$, we follow the description in P. C. Peters & J. Mathews (1963),P. C. Peters (1964), and L. D. Landau & E. M. Lifschits (1975) for the simplest case of point masses and the GW mass quadrupole formula. The length and frequency scales are normalized using a characteristic gravitational radius [^4] This corresponds to the horizon of the maximally spinning Kerr black hole with the total mass of the system., $R_{g}=G(M_{1}+M_{2})/c^{2}$ where $G$ is Newton’s gravitational constant, and $c$ is the speed of light, and the Keplerian frequency at $R_{g}$ is $\Omega_{g}=[G(M_{1}+M_{2})/R_{g}^{3}]^{1/2}$, respectively. The separation $d$ between the two NSs is expressed in terms of the orbital angular frequency $\Omega$ as
@@
where $\Omega_{i}$ is the initial orbital frequency, and
@@
This formula provides only a leading-order approximation of the rate of the energy loss from the system. For a more precise modeling, particularly during the later stages of the inspiral, higher-order corrections including tidal effects from the post-Newtonian contributions of Einstein’s equations should be employed (J. C. Lombardi et al., 1997; L. Blanchet, 2014; T. Narikawa et al., 2020; S. Yin et al., 2025). However, since the main focus of this study is on the magnetosphere structure of the BNSs and associated EM phenomena, we adopt Eq. (5) to simplify the treatment of the orbital evolution.
In our simulations, the two NSs are treated as solid conducting bodies occupying finite and fixed volumes throughout the entire evolution. Consequently, the inspiral is considered to end when the surfaces of the two stars touch, which occurs at an earlier time, $t_{\rm Final}$, than the formal merger time predicted for point masses from the GW mass quadrupole formula. This corresponds to a separation of $d_{\rm Final}=2r_{\star}$, where $r_{\star}$ is the radius of each NS. The time $t_{\rm Final}$ reads (M. V. Medvedev & A. Loeb, 2013)
@@
where, for the adopted values $r_{\star}=12\rm\,km$ and $M_{1}=M_{2}=1.4M_{\odot}$, this corresponds to a duration of $t_{0}~-~t_{\rm Final}~\sim~1.2\rm\,ms$. Hereafter, we define $\tau=t-t_{0}$ as the simulation time indicator, which is always negative, representing the time remaining until the merger. Although $\tau$ increases algebraically as the system evolves toward merger, its absolute value decreases over time.
@@ 2.4 Magnetospheric Evolution and Numerical Implementation
As the conductivity $\sigma$ varies from 0 to $\infty$, the magnetosphere structure spans the entire spectrum of solutions, from a vacuum to FF. For practical purposes, we adopt a relatively high value of conductivity, $\sigma=6\Omega_{\rm i}$, which brings the system close to the FF regime. At the end of each time step, any remaining parallel electric field component is subtracted to further enforce the ideal FF condition.
During the evolution, when using the current prescription of Eq. 10, there is a possibility for regions to develop where $E>B$ (A. Spitkovsky, 2006; C. Kalapotharakos & I. Contopoulos, 2009). Such regions are inconsistent with the FF regime due to the presence of accelerating electric field components. In those regions, the FF approximation breaks down, resulting in a superluminal plasma drift velocity, and the underlying equations become mathematically ill posed. To mitigate this issue, following A. Gruzinov (2007, 2008) and J. Li et al. (2012), we modify the Ohm’s law by including the $E_{0}^{2}$ in the denominator of the $\mathbf{E}\times\mathbf{B}$ term, as shown in Eq. 11. With this prescription, the denominator is not allowed to attain small values, facilitating the treatment of the current sheet region (see, e.g., C. Kalapotharakos et al., 2012, 2014), where the magnetic field may vanish. In real physical conditions, the plasma would be redistributed in regions where $E$ exceeds $B$ to neutralize the corresponding accelerating electric field component. Additionally, in order to prevent the development of $E>B$ regions, we enforce the condition $E\leq B$ throughout the simulation domain at the end of each time step. We do so by reducing $|\textbf{E}|$ in these regions to $98\%$ of $|\textbf{B}|$.
Our FDTD code integrates the Maxwell equations, Equations 89, tracking the inspiral of the two NSs by enforcing field values inside them that correspond to their respective magnetic moments, with the electric fields determined by Eq. (3). Compared to simulations of an isolated spinning NS, the merging BNS system proved more susceptible to the growth of oscillations (spurious ripples) near regions with steep gradients or discontinuities, such as current sheets. To address this, we implemented a third-order total variation diminishing (TVD) Runge-Kutta scheme (C.-W. Shu & S. Osher, 1988), replacing the standard third-order Runge-Kutta method used previously. This TVD scheme is well known for its enhanced numerical stability, particularly in handling sharp transitions and avoiding oscillatory artifacts.
In the case of perfect conductors, as the stars are modeled here, the boundary conditions require the continuity of the tangential components of the electric field and the normal components of the magnetic field. Handling spherical boundaries within a Cartesian grid poses significant computational challenges. This issue arises even in the simpler case of an isolated spinning NS. However, in the BNS system, the motion of the stellar boundaries aggravates the problem, often spouting unphysical nonzero values of $\nabla\cdot\mathbf{B}$ near the stellar boundary as the system evolves.
@@
Figure 3: The total Poynting flux that crosses a surface of radius $13\,r_{\star}$ during system evolution starting from the initial separation of $3.3\,r_{\star}$ (red line) and $8\,r_{\star}$ (black line). Both simulations show similar flux evolution in the case of $\alpha_{1}=0^{\circ},\,\alpha_{2}=45^{\circ}$, and $\lambda=-1$ (
@@
Various techniques have been developed to address numerical errors associated with the magnetic field divergence in MHD simulations. These include the projection method (J. Brackbill & D. Barnes, 1980), diffusive method (B. Marder, 1987; B. van der Holst & R. Keppens, 2007), constrained transport method (C. R. Evans & J. F. Hawley, 1988), eight-wave formulation approach (K. G. Powell et al., 1993), and the hyperbolic divergence cleaning approach (A. Dedner et al., 2002); see also M. Zhang & X. Feng (2016) for a comparative review.
In this work, we adopt the diffusive method scheme implemented in M. Zhang & X. Feng (2016), which adds an additional term to Eq. 9 to eliminate any residual divergence of the magnetic field. In practice, at the end of each time step, we apply the following scheme
@@
The parameter $\eta$ has to satisfy (see M. Zhang & X. Feng, 2016, and references therein)
@@
The evolution of the binary system is tracked from an initial separation, $d=3.3r_{\star}$, until the stars contact, which from Equation 7 corresponds to $t_{\rm Final}=7.7\rm\,ms$ with a time step of $dt=0.04\,r_{\star}/c$. In our simulations, we have selected a mass-radius ratio, and that defines a specific orbital evolution following the prescription described earlier  [^5] Note that the initial separation is given in units of $r_{\star}$.. Although the results presented below can be scaled with the stellar radius $r_{\star}$, in order for our simulations to still be able to describe the derived EM fluxes, for a different radius, the mass should be changed accordingly, resulting to the same mass/radius ratio. Below, if not stated otherwise, $r_{\star}=12{\rm km}$ is used, from $\Omega_{\rm i}\sim 2451\rm\,\,rad/s\,\,(390\,\rm Hz)$ to $\Omega_{\rm Final}\sim 5195\rm\,\,rad/s\,\,(827\,\rm Hz)$. The grid consists of $512$ grid points in each direction, covering the entire computational domain with an edge length of $25.6r_{\star}$ centered on the binary system. However, the active computational domain, excluding the PML region, has an edge length of $22.3r_{\star}$. This region, in practice, is where the simulation’s magnetospheric evolution is calculated.
Resolution studies were also conducted (see Appendix A). For a selected subset of configurations, we performed simulations at higher spatial resolution and smaller time steps, utilizing 1024 grid points per dimension, effectively doubling the resolution, or with computational domains twice the size of the standard setup. The results presented in §3 were found to be robust, showing no significant variation with enhanced resolution and box size.
@@ 3 Results
A series of simulations were run to survey different unique initial configurations for the magnetic fields of the two stars. We explored $\lambda=\{\pm 1,\pm 10\}$, $\alpha_{1},\alpha_{2}=\{0^{\circ},45^{\circ},90^{\circ}\}$ and $\Delta\Phi=\{0^{\circ},45^{\circ},90^{\circ}\}$. Additionally, as changes in the $\Delta\Phi$ parameter only slightly affect the outflow patterns, for $\Delta\Phi=0$, further cases were explored with $\lambda=\{\pm\sqrt{10},\pm 100\}$ and the same range for the inclination angles. Lastly, to validate the specific results presented below (see section  3.1.5), we also performed additional simulations for cases with $\lambda=\{\pm\sqrt{10^{3}},\pm\sqrt{10^{5}},\pm 10^{3}\}$, $\alpha_{1}=\{0^{\circ},45^{\circ}\}$, $\alpha_{2}=0$, and $\Delta\Phi=0$. This exploration resulted in $132=(3^{3}+3^{2}+3)2^{2}-24$ unique [^6] When $\alpha_{2}=0$, the configuration is the same for every $\Delta\Phi$; thus, those cases have been excluded. simulations.
@@
, left two columns) and antialigned (
@@
, right two columns) magnetic moments on the zxz-x plane in the corotating frame. Black streamlines indicate the magnetic field direction on the poloidal plane, while the color denotes the normalized magnitude of the magnetic field component perpendicular to it (first and third column) and the magnitude of the magnetic field (second and fourth column). Dashed vertical lines indicate the characteristic distance $c/\Omega$.
@@ 3.1 Electromagnetic Fluxes, Field Structures, and Emission Anisotropy
For a single pulsar with a plasma-filled magnetosphere, as it rotates around its spin axis with an angular frequency $\omega$, energy is lost to low-frequency EM radiation that is mostly sourced from the open magnetic field lines (I. Contopoulos et al., 1999; N. Bucciantini et al., 2006; A. Spitkovsky, 2006; C. Kalapotharakos & I. Contopoulos, 2009; A. Tchekhovskoy et al., 2016). Calculating the magnetic flux along the open magnetic field lines $\Phi_{\rm open}$, one may compute an approximate luminosity as $L\approx\omega^{2}\Phi_{\rm open}^{2}/(6\pi^{2}c)$ for direct comparison with the integrated Poynting flux crossing a surface at a large distance.
Figure 5: 3D representation of the magnetic field lines for different magnetic moment configurations. Top row: both stars have aligned magnetic moments ($\uparrow\uparrow$), shown from two perspectives, equatorial plane (left) and top-down along the orbital axis (right). Middle row: a configuration with $\alpha_{1}=0^{\circ},\alpha_{2}=45^{\circ},\Delta\Phi=0^{\circ}$ and $\lambda=-1$ (
@@
), presented at two snapshots separated by half orbital rotation. Bottom row: A configuration $\alpha_{1}=90^{\circ},\alpha_{2}=90^{\circ},\Delta\Phi=0^{\circ}$, and $\lambda=-1$ (
@@
In particular, the outgoing EM power can be calculated by integrating the radial component of the Poynting vector, $\mathbf{S}=c\mathbf{E}\times\mathbf{B}/(4\pi)$, over a spherical surface of radius $r_{F}$. Numerical FF simulations (A. Spitkovsky, 2006), later validated by MHD and particle-in-cell (PIC) simulations within the FF regime, have shown that the luminosity of a single isolated pulsar with a magnetic inclination angle $\alpha$ follows an approximate relation
@@
The Poynting flux distribution crossing a spherical surface at a given time (Fig. 2, left panel) rotates in sync with the pulsar’s rotation. By integrating this luminosity over a full rotation period, the average luminosity distribution can be determined (Fig. 2, right). An observer at infinity may be pointed at a specific set of the azimuthal angle $\phi$ and polar angle $\zeta$, i.e., the so-called observer angle relative to the spin axis. Notably, the average luminosity observed over a complete rotation depends exclusively on $\zeta$.
@@
Figure 6: The evolution of the Poynting flux, normalized to $L_{0}$ (as indicated), crossing a spherical surface of radius $13r_{\star}$ around the system for equal aligned ($\uparrow\uparrow$, left column) and antialigned ($\uparrow\downarrow$, right column) magnetic moments. For the aligned case, $L_{0}^{\uparrow\uparrow}=0.36\times 10^{44}B_{12}^{2}r_{12\,\rm km}^{2}\,\rm erg\,s^{-1}\,sr^{-1}$, while, for the antialigned case, $L_{0}^{\uparrow\downarrow}=1.24\times 10^{44}B_{12}^{2}r_{12\,\rm km}^{2}\,\rm erg\,s^{-1}\,sr^{-1}$. Here, $B_{12}$ and $r_{12\,\rm km}$ are the surface magnetic field (in units of $10^{12}\rm\,G$) and the stellar radius (in units of $1.2\times 10^{6}\rm\,cm$).
The binary system, starting from a large initial separation, evolves in a low orbital frequency regime as dictated by Eq. (4), resulting in a gradual evolution toward merger. For a synchronized system (where the stars have the same spin and orbital frequencies) in this regime, the light cylinder is mathematically well defined at a distance $R_{\rm LC}=c/\Omega$. In this configuration, the magnetosphere structure corotates with the system (neglecting the radial inspiral motion), and the Poynting flux distribution forms a fixed, rotating pattern. However, in this study, the stars are effectively irrotational, and no strictly defined single light cylinder exists. Instead, the length scale $c/\Omega$ can be interpreted as an instantaneous light cylinder associated with the evolving field configuration and orbital frequency $\Omega$. Yet, $c/\Omega$ is a crucial length scale for the magnetosphere’s evolution, and the active computational domain size is chosen so that this length scale is always well within it.
For the initial orbital frequency $\Omega_{i}$, we have $c/\Omega_{i}=10.2r_{\star}$, while, at the final time $t=t_{\text{final}}$, it decreases to $c/\Omega_{f}=4.8r_{\star}$ owing to the inspiral.
@@
Figure 7: The total Poynting flux crossing a spherical surface of radius $13r_{\star}$ throughout the system evolution for equal antialigned ($\uparrow\downarrow$, solid line) and aligned ($\uparrow\uparrow$, dashed line) magnetic moments.
3.1.1 Aligned and Antialigned Magnetic Moments
In Figure 4, we present slices of the magnetosphere’s magnetic field structure at the indicated snapshots. This comparison involves cases where the magnetic moments of both stars are either aligned ($\uparrow\uparrow$, left two columns) or antialigned ($\uparrow\downarrow$, right two columns). In Fig. 4, the slices are taken in the corotating poloidal plane, which follows the binary’s motion and contains the centers of both stars. Arrowed lines represent the corresponding poloidal magnetic field lines, and the color scale indicates the magnitude of the magnetic field component perpendicular to the poloidal plane in the first and third columns, while, in the second and fourth columns, it indicates the total magnetic field strength on a logarithmic scale.
The primary distinction between these two cases lies in the behavior of the magnetic field lines in the region around the two stars. If each star were isolated, these magnetic field lines would belong to either the closed or open magnetic field regions of their respective magnetospheres. In the presence of both stars, these lines do not maintain the same behavior. In the antialigned case, some magnetic field lines connect the two stars directly. In contrast, in the aligned case, these field lines are bent out of the poloidal plane in the azimuthal direction, closing within the opposite hemisphere of the star they originate from. This behavior is more evident in a 3D representation of the magnetic field lines, as shown in the top row of Figure 5. The left column shows the magnetic field lines from the equatorial plane, while the right column provides a top-down view. Due to the symmetry, the field lines of each star would have the same directions, and thus, no line is allowed to connect the two stars (C. Palenzuela et al., 2013a; J. F. Mahlmann & A. M. Beloborodov, 2025). Furthermore, as each star’s magnetic field lines are repelled from the region dominated by the other star’s magnetosphere, they become compressed in the leading direction of motion. In contrast, comparatively fewer field lines populate the trailing side toward the companion. Due to retardation effects, the field lines are consistently swept back, resulting in an “S”-shaped morphology when viewed from above, i.e., along the $z$-axis (see left panel of the top row in Fig. 5). The motion of the stars also results in a constant opening and closing of field lines, producing a stable flow of EM flux along the open field lines that increases following the increase of the fields as the separation decreases.